# https://github.com/beichenzbc/Long-CLIP

### v1

@@ 1 Introduction
To this end, we propose Long-CLIP as a plug-and-play alternative to CLIP. After efficiently fine-tuning CLIP using just an extra 1 million long text-image pairs with around 12 hours on 8 GPUs, Long-CLIP seamlessly supports long-text input and retains or even surpasses the zero-shot generalizability of CLIP on various benchmarks. Notably, Long-CLIP aligns the CLIP latent space, readily replacing CLIP without any further adaptation in downstream frameworks. To be specific, Long-CLIP introduces two novel designs to achieve these targets, including (1) a knowledge-preserved stretching of positional embedding and (2) a primary component matching of CLIP features.
@@ 3.1 Exploring the Effective Length of CLIP
Based on this assumption, we conduct an experiment using urban dataset, a long-caption image-text evaluation dataset constructed by us which will be further elaborated in Sec. 4.2. In theory, the R@1 should also gradually increase as we incrementally increase the input caption length if the model can leverage the additional information. However, our observations reveal that once the input length surpasses 20 tokens, the R@1 of CLIP exhibits slow growth. This suggests that the true effective length of CLIP is no more than 20 tokens, as it struggles to leverage the additional information beyond this point. In contrast, our model’s performance continues to improve as the input length increases, reaching its peak at the maximum value within the dataset. This indicates that our model is capable of consistently utilizing the newly added information in the captions.
@@ 4.1 Experiment Setting
Training setting.  We use the ShareGPT4V [2] dataset as training dataset, which contains about 1M (long caption, image) pairs. The random 1k data is separated as an evaluation dataset. We fine-tune the CLIP model for 6 epochs with batch size 2048. Detailed hyper-parameter settings will be listed in supplementary material.
@@ 4.2 New Evaluation Dataset on Long Caption Text-Image Retrieval
Table 1: The R@1 of long-caption text-image retrieval on 1k ShareGPT4V [2] validation set and 200 Urban dataset. Best result is in bold.
ShareGPT4V Urban
@@
Long-CLIP(Ours) 57.2 80.8 87.8 40.4 65.9 75.7 47.2 71.5 80.0 33.1 55.6 64.9
@@
Long-CLIP(Ours) 58.3 81.4 88.2 45.1 70.4 79.3 48.8 73.2 81.6 39.4 62.6 71.4
@@
Long-CLIP(Ours) 67.1 42.9 61.3 90.7 69.5 66.24
@@
Long-CLIP(Ours) 73.6 33.4 68.4 95.2 78.4 69.80
@@ 4.4 Ablation Study
Strategy  ImageNet  Cifar100  COCO T2I R@5  Flickr I2T R@5  urban T2I R@1
Undistinguished 65.5 67.5 64.6 66.8 71.2
@@ 5 Conclusion
Yuhang ZangJiaqi WangCorresponding author.  § Work done during an internship in Shanghai AI Laboratory

### v2

@@ 1 Introduction
To this end, we propose Long-CLIP as a plug-and-play alternative to CLIP. After efficiently fine-tuning CLIP using just an extra 1 million long text-image pairs with only 0.25 hours on 8 GPUs, Long-CLIP seamlessly supports long-text input and retains or even surpasses the zero-shot generalizability of CLIP on various benchmarks. Notably, Long-CLIP aligns the CLIP latent space, readily replacing CLIP without any further adaptation in downstream frameworks. To be specific, Long-CLIP introduces two novel designs to achieve these targets, including (1) a knowledge-preserved stretching of positional embedding and (2) a primary component matching of CLIP features.
@@ 3.1 Exploring the Effective Length of CLIP
Based on this assumption, we conduct an experiment using urban-200 dataset, a long-caption image-text evaluation dataset constructed by us, which will be elaborated in Sec. 4.2. In theory, the R@1 should also gradually increase as we incrementally increase the input caption length if the model can leverage the additional information. However, our observations reveal that once the input length surpasses 20 tokens, the R@1 of CLIP exhibits slow growth. This suggests that the true effective length of CLIP is no more than 20 tokens, as it struggles to leverage the additional information beyond this point. In contrast, our model’s performance continues to improve as the input length increases, reaching its peak at the maximum value within the dataset. This indicates that our model is capable of consistently utilizing the newly added information in the captions.
@@ 4.1 Experiment Setting
Training setting.  We use the ShareGPT4V [2] dataset as training dataset, which contains about 1M (long caption, image) pairs. The random 1k data is separated as an evaluation dataset. We fine-tune the CLIP model for 1 epoch with batch size 2048. Detailed hyper-parameter settings will be listed in supplementary material.
@@ 4.2 New Evaluation Dataset on Long Caption Text-Image Retrieval
After the first submission, we further scaled up Urban-200 into Urban-1k. The dataset has been released at https://huggingface.co/datasets/BeichenZhang/Urban1k. Urban-200 is used in the main paper. Detailed results about Urban-1k is shown in Sec. 6 in supplementary materials.
Table 1: The R@1 of long-caption text-image retrieval on 1k ShareGPT4V [2] validation set and Urban-200 dataset. Best result is in bold.
ShareGPT4V Urban-200
@@
Long-CLIP(Ours) 57.6 81.1 87.8 40.4 65.8 75.2 46.8 71.4 79.8 34.1 56.3 65.7
@@
Long-CLIP(Ours) 62.8 85.1 91.2 46.3 70.8 79.8 53.4 77.5 85.3 41.2 64.1 72.6
@@
Long-CLIP(Ours) 66.8 42.7 61.2 90.7 69.3 66.14
@@
Long-CLIP(Ours) 73.5 33.7 67.9 95.3 78.5 69.78
@@ 4.4 Ablation Study
Strategy  ImageNet  Cifar100  COCO T2I R@5  Flickr I2T R@5  Urban-200 T2I R@1
Undistinguished 65.5 67.5 64.6 66.8 71.0
@@ 5 Conclusion
Yuhang ZangJiaqi WangCorresponding author.  § Work done during an internship in Shanghai AI Laboratory. This work is partially supported by the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial lntelligence Laboratory.
6 Urban-1k Dataset
Urban-1k is a scaling-up version of Urban-200 dataset in the paper. It contains 1k urban images and their corresponding captions generated by GPT-4V. Each caption contains about 107 words on average. The process for building Urban-1k dataset is exactly the same as Urban-200. It has been released at https://huggingface.co/datasets/BeichenZhang/Urban1k.
Table 6: The result on Urban-1k dataset. Best result is in bold.
Model  Urban-1k I2T R@1  Urban-1k T2I R@1
B/16 CLIP 68.1 53.6
Long-CLIP(Ours) 78.9 79.5
L/14 CLIP 68.7 52.8
Long-CLIP(Ours) 82.7 86.1